OpenStudy (anonymous):

A rare bacterial culture is being grown in a lab. As the days progress, the cells multiply and grow. After 1 day, there is only 1 cell. After 2 days, there are 9 cells. After 3 days, there are 20 cells. After 4 days, there are 34 cells. Which recursive equation represents the pattern?

OpenStudy (anonymous):

an = an – 1 + 3n an = an – 1 + n + 2) an = an – 1 + 3(n + 2) an = an – 1 + (3n + 2)